OpenStudy (anonymous):

X-6=square root -7x+72 sorry no radical symbol on iPad

OpenStudy (anonymous):

\[x-6=\sqrt{-7x+72}\]

OpenStudy (anonymous):

Square the whole equation and get a quadratic

Directrix (directrix):

x - 6 = sqrt(-7x + 72) Square both sides (x-6)^2 = [ sqrt(-7x + 72) ] 2 (x-6)^2 = -7x + 72 x^2 - 12x + 36 = -7x + 72 x^2 -5x - 36 = 0 (x - 9) ( x + 4) = 0 x = 9 or x = -4 x = 9 Note: x = -4 is discarded because the principal square root of an expression is never equal to a negative. If x were -4, then x - 6 = sqrt(-7x + 72) would begin as -4 -6 = sqrt(....) which cannot happen in the set of Real numbers.